OpenStudy (anonymous):

need some math help please asap Gino designed a garden for his rectangular backyard. The design he drew on paper is ten inches long. The length of his backyard is forty feet. If his drawing is to scale, and his backyard is twenty-eight feet wide, how many inches wide is his drawing? Write a proportion using words. Then, use the proportion to calculate the drawing width. Show your work.

OpenStudy (anonymous):

10 inches = 40 feet divide both sides by 10 1 inch = 4 feet 28 feet / 4 feet = 7 so it's 7 inches

OpenStudy (anonymous):

since the drawing can be scaled and the ratio of a feature from drawing:garden is coonstant, we can write, constant scale ratio=10/40=1/4=drawing length/actual length similarly writing drawing width/actual width=1/4 and so we get drawing width=7

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

is this answer right?